Find the value of a for which the polynomial x^3- a x^2- ax +1 has −1 as a root with multiplicity at least 2.
1
Expert's answer
2020-02-24T10:26:58-0500

We know that:

If "x = a" is a root of a polynomial then "(x - a)" will be the factor of this polynomial.

If the multiplicity of this root "a" is "n" of the polynomial then "(x-a)^n" will be the factor of the polynomial.

Actually, the multiplicity of the root is the power value of a factor ( Here, This root is involved in this factor) of the polynomial.

If the polynomial is the degree of "m" then the highest power of the variable in the polynomial expression is "m"

And:

If the polynomial has a degree "m" then this polynomial is created by using "m" number of factors.

Now:

The polynomial is given by:

"x^3 - ax^2 -ax + 1"

Since:

The root of the given polynomial is "-1"

So:

"(x - (-1)) \, \, or \,\, (x+1)" will be one factor of the given polynomial.

Again:

Since the root "-1" has multiplicity at least "2"

So:

"(x + 1)^2" will be the factors of the given polynomial.

Now:

We can see that our given polynomial has degree three so we have to required three factors but we have two factors which are "(x+1)^2"

For the third factors, we will assume it. As the last term of our polynomial is "1" so our final factor must be "(x + 1)" otherwise we can not get the last term "1" of the given polynomial.

Hence:

We can write:

"x^3 - ax^2 - ax + 1 \,\,\, = (x+1)^3 \\
\hspace{1 cm} \hspace{1 cm} \hspace{1 cm} = x^3 + 3x^2 + 3x + 1"

Comparing both of side, we have:

"a = -3"

[ From the given information, it may also prove that the power of "( x +1)" must be "3" because the root "x = -1" of the given polynomial has a multiplicity which is not exactly "2" but at least "2" ]